OpenStudy (anonymous):

What is the solution to the equation 9^-3x ≈ 7 ?

OpenStudy (anonymous):

are you there?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

what class is this for? if any

OpenStudy (anonymous):

Algebra II

OpenStudy (anonymous):

are you doing logarithms?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

is it multiple choice?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

x=\[\frac{ \log(7) }{ 6 \log(3) }\] is this one of your answer choices?

OpenStudy (anonymous):

x = -0.295 x = -0.376 x = 0.376 x = 0.295

hero (hero):

@iplayffxiv, what kind of website do you think this is? A site where you help students understand how to do the problem or a site where you just give the answer?

OpenStudy (anonymous):

Do you know how to do this, Hero?

OpenStudy (anonymous):

uhhh a site where u help people understand............

hero (hero):

@celecity, yes, I know how to do it. @iplayffxiv, how should I classify what you have done here? Giving help or giving answers?

OpenStudy (anonymous):

Would you mind walking me through it?

OpenStudy (anonymous):

uhhhhh i dont know.............

OpenStudy (anonymous):

why dont you tell me

hero (hero):

Well, @iplayffxiv, I don't see where you explained any concepts or gave any general formulas, or defined anything or presented any postulates or anything related to general problem solving. All I see that you've posted is a possible answer choice.

hero (hero):

Anyway, @celecity, \[9^{-3x} \approx 7\]

hero (hero):

Or more explicitly, \[9^{(-3x)} \approx 7\]

hero (hero):

Anytime you see a negative exponent, it translates to the multiplicative inverse of the given expression. Therefore we can re-write the expression as the inverse of \(9^{3x}\). Doing that will give us: \[\frac{1}{9^{3x}} \approx 7\]

hero (hero):

Now, there's a particular rule we can use to simplify the expression in the denominator: \(a^{(bc)} = (a^{b})^c\)

hero (hero):

So \(9^{3x} = (9^3)^x\) or simply \(729^x\)

hero (hero):

Replacing that in the denominator, we have: \[\frac{1}{729^x} = 7\]

hero (hero):

Now, since we are trying to isolate x, let's write this in a more convenient manner. Muliply both sides by \(729^x\) and divide by sides by 7 to get: \[\frac{1}{7} = 729^x\] Do you think you might be able to solve it from here?

hero (hero):

Hint: Taking logs of both sides is involved.

OpenStudy (anonymous):

So that would be log1/7=log729?

hero (hero):

What did you do with the x?

hero (hero):

Remember, there was an x in the exponent before you logged both sides. What did you do with it?

OpenStudy (anonymous):

It would be =log729^x

OpenStudy (anonymous):

Right?

hero (hero):

Then after logging both sides, what will the resultant equation be?

OpenStudy (anonymous):

The log1/7 results in -0.84509804 and the log729^x is 28.62727528

hero (hero):

Remember \(\log a^x = x \log(a)\)

OpenStudy (anonymous):

Oh, okay, one moment.

OpenStudy (anonymous):

-0.84509804 = x(2.862727528)

hero (hero):

I think you might know what the next step is in order to isolate x. What does x equal?

OpenStudy (anonymous):

-0.2952072916

hero (hero):

Is that one of your answer choice?

OpenStudy (anonymous):

-0.295 is. Thank you so much for your help :)

hero (hero):

yw